OpenStudy (anonymous):

find the exact value of cos-1 (square root of 3 divided by 2)

OpenStudy (anonymous):

Looks like this? \[\cos^{-1}(\sqrt{\frac{3}{2}})\]

OpenStudy (anonymous):

Thats right

OpenStudy (lalaly):

30 degrees, pi/6 in radians

OpenStudy (anonymous):

thanks :)

OpenStudy (lalaly):

:D:D

OpenStudy (anonymous):

Ah, must have been \[\cos^{-1}(\frac{\sqrt{3}}{2})\] Inverse cosine isn't defined at the previously typed expression.

OpenStudy (anonymous):

Oh. shoot. lol